OpenStudy (anonymous):

For a holiday party, Kathleen is making a six-layer torte, which consists of six 8-inch diameter cake layers with a cream topping on each layer. She knows that one recipe of cream will cover 50 square inches of cake. How many batches of the topping must she make to complete her torte? ( A = p )(Use p = 3.14 as an approximate value)

OpenStudy (anonymous):

6 layers = 6 layers of cream 8 in diameter -> 4 in radius \[AreaOfALayer = r^2 π \] \[BatchOfCream = 50 in^2\] \[Batches*50*in^2 = (6*layers)( 4^2 π*in^2/layer)\] After solving for the number of batches I get 6.03 batches. Thus, to make sure that there is enough cream, 7 batches of cream should be made. Work attached.